OpenStudy (anonymous):

coniser a point charge q=1 mc placed at corner of cube of sides 10 cm find the flux through each face of cube

OpenStudy (sidsiddhartha):

here the charge is placed in the corner of the cube so in 3D u can see the charge is enclosed by eight cubes so according to gauss law flux through surfaces will be=q/(8*e) here q=1mc,e=8.85*10^-12